### v2

# Collective order boundedness of sets of operators between ordered vector spaces

Abstract
It is proved that: each collectively order continuous set of operators between Archimedean ordered vector spaces with generating cones is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded. Several applications to commutative operator semigroups on ordered vector spaces are given.
Keywords: ordered vector space, collectively qualified set of operators, commutative operator semigroup
MSC2020: 46A40, 46B42, 47B60, 47D03
1 Introduction
Collectively compact sets of operators between normed spaces were introduced by P. M. Anselone and T. W. Palmer [5]. Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 7, 8, 9]. The present paper generalizes results of [1, 7, 15] to the ordered vector space setting. The results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator between Archimedean ordered vector spaces is order bounded; 2) each order to norm continuous operator from an ordered Banach space with a closed generating cone to a normed space is continuous.
We shall abbreviate a normed space (ordered vector space, ordered normed space, ordered Banach space, vector lattice, normed lattice, Banach lattice) by NS (OVS, ONS, OBS, VL, NL, BL). Throughout the text, vector spaces are real, operators are linear, symbol ${\cal L}(X,Y)$ (${\cal L}_{+}(X,Y)$, ${\cal L}_{ob}(X,Y)$, ${\cal L}_{oc}(X,Y)$, ${\cal L}_{onb}(X,Y)$) stands for the space of all (positive, order bounded, order continuous, order to norm bounded) operators from $X$ to $Y$. A symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
Quite often, working with collective properties of a set of operators in OVSs involves order convergent nets. Let $X$ be an OVS. A net $(x_{\alpha})$ in $X$ is
  1. -
    order convergent to $x\in X$ (briefly, o-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) whenever there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is an $\alpha_{\beta}$ satisfying $-g_{\beta}\leq x_{\alpha}-x\leq g_{\beta}$ for all $\alpha\geq\alpha_{\beta}$.
  2. -
    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
We shall use the following OVS version of [7, Definition 2.1]. Let $X$ be an OVS, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$. We say that ${\cal B}$
  1. -
    collective o-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-o}}0$) if there exists a net $g_{\beta}\downarrow 0$ such that, for each $\beta$, there is an $\alpha_{\beta}$ satisfying ±$\pm x^{b}_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$ and $b\in B$.
  2. -
    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
[…]
The following definition extends [7, Definition 2.1] to OVSs.
Definition 1.1.
Let ${\cal T}\subseteq{\cal L}(X,Y)$.
When $X$ and $Y$ are OVSs,
  1. -
    ${\cal T}$ is collectively order bounded (briefly, ${\cal T}\in{\bf L}_{ob}(X,Y)$) if, for every order interval $[a,b]$ in $X$ there exists an order interval $[u,v]$ in $Y$ such that ${\cal T}[a,b]:=\bigcup\limits_{T\in\mathcal{T}}{T}[a,b]\subseteq[u,v]$.
  2. -
    ${\cal T}$ is collectively o-continuous (briefly, ${\cal T}\in\text{\bf L}_{oc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-o}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$;
  3. -
    ${\cal T}$ is collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
When $X$ is an OVS and $Y$ is an ONS,
  1. -
    ${\cal T}$ is collectively almost order bounded (briefly, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$) if, for every $[a,b]\subseteq X$ and $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}[a,b]\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is a normed space and $Y$ is an ONS,
  1. -
    ${\cal T}$ is collectively semicompact (briefly, ${\cal T}\in\text{\bf L}_{sc}(X,Y)$) if, for every $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}B_{X}\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is an OVS and $Y$ is a normed space,
  1. -
    ${\cal T}$ is collectively order to norm bounded (briefly, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$) if the set ${\cal T}[a,b]$ is norm bounded for every $[a,b]\subseteq X$.
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an OVS to an Archimedean OVS is collectively order bounded. Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets. Finally, we prove Theorem 2.3 saying that under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded. It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded. Section 3 is devoted to applications of obtained results to commutative operator semigroups on OVSs.
For the terminology and notations not explained in the text, we refer to [2, 3, 6, 13, 14].
2 Main results
We are going to prove the following results which generalize the results obtained recently in [7, 15] in the vector lattice setting to the OVS setting.
  1. -
    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS with a generating cone is collectively order bounded.
  1. -
    Every collectively order to norm bounded set of operators from an OBS with a closed generating cone to a NS is norm bounded
  1. -
    Collectively r-continuous sets coincides with collectively order bounded sets when the OVS in the range has a generating cone.
2.1
We begin with generalization of the first part of [7, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
Theorem 2.1.
Let $X$ and $Y$ be OVSs with generating cones and suppose $X$ is Archimedean. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
Proof.
Let ${\cal T}\in\text{\bf L}_{oc}(X,Y)$. It is enough to prove order boundedness of ${\cal T}[0,b]$ for every $b\in X_{+}$. Take $[0,b]\subseteq X$. Following [1, Theorem 2.1], consider the lexicographically directed set =×${\cal I}=\mathbb{N}\times[0,b]$: $(m,z)\geq(n,z^{\prime})$ if $m>n$ or $m=n$ and $z\geq z^{\prime}$, and let $x_{(k,z)}=\frac{1}{k}z$. Since $X$ is Archimedean, $\frac{1}{n}b\downarrow 0$. It follows from $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$ for $(m,z)\geq(n,b)$ that $x_{(m,z)}\operatorname{\xrightarrow[]{o}}0$.
Since ${\cal T}\in\text{\bf L}_{oc}(X,Y)$, there exists $g_{\beta}\downarrow 0$ in $Y$ such that, for every $\beta$ there exists $(m_{\beta},y_{\beta})$ satisfying ±$\pm Tx_{(m,y)}\leq g_{\beta}$ for all $T\in{\cal T}$ and $(m,y)\geq(m_{\beta},y_{\beta})$. Pick any $g_{\beta_{0}}$. Since $(m_{\beta_{0}}+1,y)\geq(m_{\beta_{0}},y_{\beta_{0}})$,
±$\pm T\Big{(}\frac{y}{m_{\beta_{0}}+1}\Big{)}=\pm Tx_{(m_{\beta_{0}}+1,y)}\leq g_{\beta_{0}}\ \ \ (y\in[0,b],T\in{\cal T}).$
Therefore, ${\cal T}[0,b]\subseteq[-(m_{\beta_{0}}+1)g_{\beta_{0}},(m_{\beta_{0}}+1)g_{\beta_{0}}]$. ∎
In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] as follows.
Corollary 2.1.
Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $Y$ with generating cone is order bounded.
The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Corollary 2.1 even in the VL setting.
Example 2.1.
Let $X$ be the lexicographically ordered real plane $\mathbb{R}^{2}$, $Y$ the pointwise ordered $\mathbb{R}^{2}$, and $T:X\to Y$ the identity operator. Both $X$ and $Y$ are VLs. The operator $T$ is not order bounded since the order interval $[(0,0),(2,0)]$ of $X$ contains line $x=1$ that is not order bounded in $Y$. However, $T$ is order continuous. To see this, it is enough to show that $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ implies $T((x_{\alpha},y_{\alpha}))=(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$. Suppose $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ and let $\varepsilon>0$. There exists $\alpha_{0}$ with $x_{\alpha}=0$ for $\alpha\geq\alpha_{0}$. Now, take $\alpha_{1}\geq\alpha_{0}$ such that $0\leq\max\{x_{\alpha},y_{\alpha}\}=y_{\alpha}\leq\varepsilon$ for $\alpha\geq\alpha_{1}$. It follows that $T(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$.
2.2
Next, we generalize [15, Theorem 5.1] and [7, Theorem 2.2] to OVSs.
Theorem 2.2.
Let $X$ and $Y$ be OVSs with generating cones then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
Proof.
Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. In order to show ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$. Take $[0,b]\subseteq X$. Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1. If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$. Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{r}}0$ in $X$.
Since ${\cal T}\in\text{\bf L}_{rc}(X,Y)$, there exist $u\in Y_{+}$ and an increasing sequence $(k_{n},z_{n})$ in ${\cal I}$ satisfying ±$\pm T\big{(}\frac{z}{k}\big{)}=\pm Tx_{(k,z)}\leq\frac{1}{n}u$ for all $T\in{\cal T}$ and $(k,z)\geq(k_{n},z_{n})$. In particular, ±$\pm T\big{(}\frac{z}{k_{1}+1}\big{)}=\pm Tx_{(k_{1}+1,z)}\leq u$ for all $T\in{\cal T}$ and $z\in[0,b]$. So, ${\cal T}[0,b]\subseteq[-(k_{1}+1)u,(k_{1}+1)u]$. It follows ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
Now, suppose $Y=Y_{+}-Y_{+}$. Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$. Then, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $-u\leq nx_{\alpha}\leq u$ (and hence $0\leq u+nx_{\alpha}\leq 2u$) for all $\alpha\geq\alpha_{n}$. Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$, there exist $w_{1},w_{2}\in Y_{+}$ satisfying $Tu\in[-w_{1},w_{1}]$ and $T(u+nx_{\alpha})\in[-w_{2},w_{2}]$ for all $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. Then $nTx_{\alpha}\in[-(w_{1}+w_{2}),(w_{1}+w_{2})]$ (and hence ±$\pm Tx_{\alpha}\leq\frac{w_{1}+w_{2}}{n}$) for $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. It follows ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. ∎
Corollary 2.2.
Order bounded operators between OVSs with generating cones agree with r-continuous operators.
2.3
Recall that order intervals in an OBS $Y$ with a closed cone are norm bounded iff the cone is normal (cf., [3, Theorem 2.40]). We include the following elementary collective extension of this fact .
Proposition 2.1.
Let $X$ be an OVS and $Y$ be an OBS with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
Proof.
Let ${\cal T}\in\text{\bf L}_{aob}(X,Y)$ and take $[a,b]\subseteq X$. Then $\bigcup_{T\in\mathcal{T}}{T}[a,b]\subseteq[-y_{1},y_{1}]+B_{Y}$ for some $y_{1}\in Y_{+}$. Since $Y_{+}$ is normal, $[-y_{1},y_{1}]+B_{Y}$ is norm bounded by [3, Theorem 2.40]. Therefore, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. ∎
The normality of $Y_{+}$ is essential in Proposition 2.1. Indeed, the identity operator in the OBS $C^{1}[0,1]$ of continuously differentiable functions on $[0,1]$ is order bounded yet not order to norm bounded.
The following result generalizes [7, Theorem 2.1] to OBSs.
Theorem 2.3.
Let $X$ be an OBS with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
Proof.
Let ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. Suppose in contrary ${\cal T}B_{X}$ is not norm bounded. The M. G. Krein – Smulian theorem (cf., [3, Theorem 2.37]) implies $\alpha B_{X}\subseteq B_{X}\cap X_{+}-B_{X}\cap X_{+}$ for some $\alpha>0$, and hence ${\cal T}(B_{X}\cap X_{+})$ is not norm bounded. Pick sequences $(x_{n})$ in $B_{X}\cap X_{+}$ and $(T_{n})$ in ${\cal T}$ with $\|T_{n}x_{n}\|\geq n^{3}$ for all $n$, and set $x:=\|\cdot\|-$$\sum\limits_{n=1}^{\infty}n^{-2}x_{n}\in X_{+}$. Since ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ then ${\cal T}[0,x]\subseteq MB_{Y}$ for some $M<\infty$. It follows from $n^{-2}x_{n}\in[0,x]$ that $\|T_{n}(n^{-2}x_{n})\|\leq M$ for every $n$. This is absurd, because $\|T_{n}(n^{-2}x_{n})\|\geq n$ for all $n$. The proof is complete. ∎
It is worth noting that even in the case of a single operator, Theorem 2.3 generalizes the well known fact that every order bounded operator between BLs is bounded. The condition $X_{+}-X_{+}=X$ is essential in Theorem 2.3.
Example 2.2.
Take any infinite dimensional Banach space $X$ with the trivial cone $\{0\}$, consider arbitrary discontinuous operator $T$ on $X$, and observe that $T$ is order to norm bounded simply because order intervals in $X$ are singletons.
3 Some applications to operator semigroups
We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case). More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [13, p.75]).
Proposition 3.1.
Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$, and hence for all $S\in{\text{\rm co}}(\mathcal{G})$. Thus,
$\bigcup_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}SB_{X}=S_{0}\Big{(}\bigcup_{S\in{\text{\rm co}}(\mathcal{G})}SB_{X}\Big{)}\subseteq S_{0}(MB_{X})=MS_{0}B_{X}.$
Since $S_{0}B_{X}$ is relatively (weakly) compact, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. ∎
Combining Propositions 3.1 and 2.1, and keeping in mind that every OBS with a generating normal cone is Archimedean, we obtain.
Corollary 3.1.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an OBS $X$ with a closed, generating, and normal cone. If $\mathcal{G}$ contains a (weakly) compact operator $S_{0}$ then $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
Next, we consider commutative operator semigroups on AL-spaces.
Proposition 3.2.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on on an AL-space $E$ containing a weakly compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$. Then $\|S\|\leq M<\infty$ for all $S\in{\text{\rm co}}(\mathcal{G})$. We conclude $\bigcup_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}SB_{E}\subseteq MS_{0}B_{E}$. Since the set $S_{0}B_{E}$ is relatively weakly compact, it is almost order bounded by the Dunford-Pettis theorem (cf., [14][Theorem 2.5.4]). Thus, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. ∎
Combining Propositions 3.2 and 2.1 we obtain.
Corollary 3.2.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an AL-space $E$ containing a weakly compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
Finally, we include the following convergence theorem.
Theorem 3.1.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an order continuous BL $E$. Then the net $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$, and hence for all $S\in\overline{{\text{\rm co}}}(\mathcal{G})$. Let $x\in E_{+}$. By the assumption, the set $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is almost order bounded. Since $E$ is order continuous, $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is relatively weakly compact due to the Grothendieck lemma (cf., [2, Theorem 3.44]).
Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [13, Theorem 1.5]). Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
Corollary 3.3.
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a one-parameter discrete operator semigroup on an order continuous BL $E$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on OVSs.
The next two lemmas follow directly from [7, Theorems 2.1 and 2.2] and from Theorems 2.1, 2.2, and 2.3.
Lemma 3.1.
Every collectively order to norm bounded set of operators on an (OBS) with a closed generating cone is norm bounded.
Lemma 3.2.
Let $X$ be and OVS with generating cone. Then $\text{\bf L}_{rc}(X)=\text{\bf L}_{ob}(X)$. If $X$ is additionally Archimedean then $\text{\bf L}_{oc}(X)\subseteq\text{\bf L}_{ob}(X)$.
Next definition localizes several notions concerning semigrops.
Definition 3.1.
A semigroup $(T_{s})_{s\geq 0}$ on an OVS $X$ is
  1. $a)$
    locally order bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M$.
  2. $b)$
    locally o-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M$.
  3. $c)$
    locally ru-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{rc}(X)$ for every $M$.
A semigroup $(T_{s})_{s\geq 0}$ on a normed OVS $X$ is
  1. $e)$
    locally order to norm bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$
    for every $M$.
In several important cases, in each of the four items of Definition 3.1 the condition “for every $M$” can be replaced by “for some $M$”.
Proposition 3.3.
Let $(T_{s})_{s\geq 0}$ be a semigroup on an OVS $X$ such that $X_{+}-X_{+}=X$.
  1. $i)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order bounded.
  2. $ii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{rc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally ru-continuous.
  3. $iii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order continuous.
  4. $iv)$
    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order to norm bounded.
Proof.
$i)$ Let $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Let $n\varepsilon\geq M$, and take arbitrary $a\leq b$ in $X$. It follows from $X_{+}-X_{+}=X$ that $[a,b]\subseteq[-a_{0},a_{0}]$ for some $a_{0}\in X_{+}$. Since $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ and $X_{+}$ is generating, $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-a_{1},a_{1}]$ for some $a_{1}\in X$. Similarly, for some $a_{2}\in X$, we have $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}]$ and
$\bigcup\limits_{\varepsilon\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]=T_{\varepsilon}\left(\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\right)\subseteq T_{\varepsilon}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}].$
It follows $\bigcup\limits_{0\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-(a_{1}+a_{2}),a_{1}+a_{2}]$. Repeating n-times the argument, we obtain $a_{1},a_{2},...,a_{n}\in X$ with
$\bigcup\limits_{0\leq s\leq M}T_{s}[a,b]\subseteq\bigcup\limits_{0\leq s\leq n\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq\left[-\sum_{i=1}^{n}a_{i},\sum_{i=1}^{n}a_{i}\right].$ (1)1( 1 )
Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order bounded.
$ii)$ By Lemma 3.2, it follows from $i)$.
$iii)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, and fix $M\geq 0$. Let $n\varepsilon\geq M$ and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Pick a net $g^{(1)}_{{\beta_{(1)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(1)}$ there is an $\alpha_{\beta_{(1)}}$ with ±$\pm T_{s}x_{\alpha}\leq g^{(1)}_{{\beta_{(1)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(1)}}$. Since $T_{\varepsilon}$ is order continuous, $T_{\varepsilon}x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. It follows from $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ that there exists a net $g^{(2)}_{{\beta_{(2)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(2)}$ there is $\alpha_{\beta_{(2)}}$ with ±$\pm T_{s}T_{\varepsilon}x_{\alpha}\leq g^{(2)}_{{\beta_{(2)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(2)}}$. Repeating n-times, we obtain nets
$g^{(1)}_{{\beta_{(1)}}}\downarrow 0,\ g^{(2)}_{{\beta_{(2)}}}\downarrow 0,\ ...\ g^{(n)}_{{\beta_{(n)}}}\downarrow 0$
so that, for each $\beta_{(i)}$, $i=1,2,...,n$, there is $\alpha_{\beta_{(i)}}$ satisfying
±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ \alpha\geq\alpha_{\beta_{(i)}}).$
Consider the net $g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}=\sum_{i=1}^{n}g^{(i)}_{{\beta_{(i)}}}\downarrow 0$. Take an
$\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\geq\alpha_{\beta_{(1)}},\alpha_{\beta_{(2)}},....,\alpha_{\beta_{(n)}}.$
Then, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ i=1,2,...,n).$
Therefore, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
±$\pm T_{s}x_{\alpha}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq n\varepsilon).$ (2)2( 2 )
Since $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ is arbitrary, (2) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq n\varepsilon}\in\text{\bf L}_{oc}(X)$, and hence $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order continuous.
$iv)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Let $n\varepsilon\geq M$ and take some $a\in X_{+}$. By Lemma 3.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$. Then
$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C\|a\|.$ (3)3( 3 )
Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order to norm bounded. ∎
In recent works [10, 11, 12] relatively uniformly continuous semigroups on VLs were deeply studied. The use of collectively qualified semigroups is to rough for getting progress in this direction. We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.