### v4

# Collective order boundedness of sets of operators between ordered vector spaces

Abstract
It is proved that: each collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to an ordered vector space is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded. Several applications to commutative operator semigroups on ordered vector spaces are given.
Eduard Emelyanov1, Nazife Erkurşun-Özcan2, Svetlana Gorokhova 3
$1$ Sobolev Institute of Mathematics, Novosibirsk, Russia
$2$ Hacettepe University, Ankara, Turkey
$3$ Southern Mathematical Institute, Vladikavkaz, Russia
Keywords: Ordered vector space, collectively qualified set of operators, commutative operator semigroup
MSC2020: 46A40, 46B42, 47B60, 47D03
1 Introduction
Collectively compact sets of operators between normed spaces were introduced by Anselone and Palmer [5]. Recently, collectively qualified sets of operators were investigated in [4, 10, 11, 12, 14]. In the present paper, we generalize some results of [1, 8, 10, 18] to collectively qualified sets in the ordered vector space setting. Results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator on an Archimedean ordered vector space with a generating cone is order bounded; 2) each order to norm bounded operator on an ordered Banach space with a closed generating cone is continuous.
Throughout the text, vector spaces are real, operators are linear, ${\cal L}(X,Y)$ stands for the space of operators from a vector space $X$ to a vector space $Y$, and $x_{\alpha}\downarrow 0$ stands for a decreasing net in an ordered vector space such that $\inf\limits_{\alpha}x_{\alpha}=0$.
Using of order convergence is common in study of collective properties in ordered vector spaces. A net $(x_{\alpha})$ in an ordered vector space $X$ is
  1. -
    order convergent to $x\in X$ (briefly, o-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) whenever there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $-g_{\beta}\leq x_{\alpha}-x\leq g_{\beta}$ for all $\alpha\geq\alpha_{\beta}$.
  2. -
    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
We shall use the following version of [10, Definition 1.2]. Let $X$ be an ordered vector space, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$. We say that ${\cal B}$
  1. -
    collective o-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-o}}0$) if there exists a net $g_{\beta}\downarrow 0$ such that, for each $\beta$, there is $\alpha_{\beta}$ satisfying ±$\pm x^{b}_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$ and $b\in B$.
  2. -
    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-ru}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
[…]
Definition 1.1.
(cf., [10, Definition 2.1]) Let ${\cal T}\subseteq{\cal L}(X,Y)$.
When $X$ and $Y$ are ordered vector spaces, the set ${\cal T}$ is
  1. $a)$
    collectively order bounded (briefly, ${\cal T}\in{\bf L}_{ob}(X,Y)$) if, for every order interval $[a,b]$ in $X$ there exists an order interval $[u,v]$ in $Y$ such that ${\cal T}[a,b]:=\bigcup\limits_{T\in\mathcal{T}}{T}[a,b]\subseteq[u,v]$.
  2. $b)$
    collectively o-continuous (briefly, ${\cal T}\in\text{\bf L}_{oc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-o}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$;
  3. $c)$
    collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-ru}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$;
When $X$ is an ordered vector space and $Y$ an ordered normed space, the set ${\cal T}$ is
  1. $d)$
    collectively almost order bounded (briefly, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$) if, for every $[a,b]\subseteq X$ and $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}[a,b]\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is an ordered vector space and $Y$ is a normed space, the set ${\cal T}$ is
  1. $e)$
    collectively order to norm bounded (briefly, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$) if the set ${\cal T}[a,b]$ is norm bounded for every $[a,b]\subseteq X$.
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to an ordered vector space is collectively order bounded. Theorem 2.4 describes relationship between collectively ru-continuous and collectively order bounded sets, and Theorem 2.8 says that, under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded. As a consequence, Corollary 2.9 says that each order to norm bounded operator from an ordered Banach space with a closed generating cone to a normed space is continuous. Proposition 2.10 tell us that every order bounded operator between ordered Banach spaces with closed generating cones is continuous. Section 3 is devoted to applications of obtained results to operator semigroups on ordered vector spaces.
For the terminology and notations that are not explained in the text, we refer to [2, 3, 9, 17].
2 Main results
We are going to prove the following results which generalize the corresponding ones obtained in [1, 10, 18] to the setting of ordered vector spaces.
  1. -
    Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to any ordered vector space is collectively order bounded.
  1. -
    Theorem 2.4 saying that collectively ru-continuous sets of operators between ordered vector spaces with generating cones agree with collectively order bounded sets.
  1. -
    Theorem 2.8 saying that every collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space norm bounded
Let us begin with the following generalization of the first part of [10, Theorem 2.2] which is a collective version of the result by Abramovich and Sirotkin. The proof is based on ideas in the proof of [1, Theorem 2.1].
Theorem 2.1.
Let $X$ be an Archimedean ordered vector space with a generating cone. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$ for every ordered vector space $Y$.
Proof.
Let ${\cal T}\in\text{\bf L}_{oc}(X,Y)$. Since $X_{+}-X_{+}=X$, it is enough to prove order boundedness of ${\cal T}[0,b]$ for every $b\in X_{+}$. Take $[0,b]\subseteq X$. Following [1, Theorem 2.1], consider the lexicographically directed set =×${\cal I}=\mathbb{N}\times[0,b]$: $(m,z)\geq(n,z^{\prime})$ if $m>n$ or $m=n$ and $z\geq z^{\prime}$, and let $x_{(k,z)}=\frac{1}{k}z$. Since $X$ is Archimedean, $\frac{1}{n}b\downarrow 0$. It follows from $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$ for $(m,z)\geq(n,b)$ that $x_{(m,z)}\operatorname{\xrightarrow[]{o}}0$.
Since ${\cal T}\in\text{\bf L}_{oc}(X,Y)$, there exists $g_{\beta}\downarrow 0$ in $Y$ such that, for every $\beta$ there exists $(m_{\beta},y_{\beta})$ satisfying ±$\pm Tx_{(m,y)}\leq g_{\beta}$ for all $T\in{\cal T}$ and $(m,y)\geq(m_{\beta},y_{\beta})$. Pick any $g_{\beta_{0}}$. Since $(m_{\beta_{0}}+1,y)\geq(m_{\beta_{0}},y_{\beta_{0}})$,
±$\pm T\Big{(}\frac{y}{m_{\beta_{0}}+1}\Big{)}=\pm Tx_{(m_{\beta_{0}}+1,y)}\leq g_{\beta_{0}}\ \ \ (y\in[0,b],T\in{\cal T}).$
Therefore, ${\cal T}[0,b]\subseteq[-(m_{\beta_{0}}+1)g_{\beta_{0}},(m_{\beta_{0}}+1)g_{\beta_{0}}]$. ∎
In particular, this extends the Abramovich–Sirotkin theorem as follows.
Corollary 2.2.
Each order continuous operator from an Archimedean ordered vector space with generating cone to any ordered vector space is order bounded.
Note that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Theorem 2.1 and in Corollary 2.2 even in the vector lattice setting.
Example 2.3.
Let $X$ be the lexicographically ordered real plane $\mathbb{R}^{2}$, $Y$ the coordinate-wise ordered $\mathbb{R}^{2}$. Both $X$ and $Y$ are vector lattices. The identity operator $T:X\to Y$ is not order bounded since the order interval $[(0,0),(2,0)]$ of $X$ contains line $x=1$ that is not order bounded in $Y$. However, $T$ is order continuous. To see this, it is enough to show that $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ implies $T((x_{\alpha},y_{\alpha}))=(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$. Suppose $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ and let $\varepsilon>0$. There exists $\alpha_{0}$ with $x_{\alpha}=0$ for $\alpha\geq\alpha_{0}$. Now, take $\alpha_{1}\geq\alpha_{0}$ such that $0\leq\max\{x_{\alpha},y_{\alpha}\}=y_{\alpha}\leq\varepsilon$ for $\alpha\geq\alpha_{1}$. It follows that $T(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$.
An inspection of the proof of [1, Theorem 2.1] shows that Abramovich and Sirotkin proved already that each ru-continuous operator between vector lattices $E$ and $F$ is order bounded. Conversely, each order bounded operator $T:E\to F$ is ru-continuous since, for every $u\in E_{+}$ there exists $w\in F$ such that $T[-u,u]\subseteq[-w,w]$, and hence $|x_{\alpha}|\leq\frac{1}{n}u$ implies $|Tx_{\alpha}|\leq\frac{1}{n}w$. This assertion was independently proved by Taylor and Troitsky in [18, Theorem 5.1]. Note that this fact can also be derived from the nonstandard criteria of order boundedness of operators [8, Theorem 1.7.2] obtained by the first author. The following theorem generalizes this fact to the setting of ordered vector space.
Theorem 2.4.
Let $X$ and $Y$ be ordered vector spaces.
  1. $i)$
    If $X$ has a generating cone then $\text{\bf L}_{ruc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
  2. $ii)$
    If $Y$ has a generating cone then $\text{\bf L}_{ob}(X,Y)\subseteq\text{\bf L}_{ruc}(X,Y)$.
Proof.
$i)$   Let ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$. Since $X_{+}-X_{+}=X$, to show ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$. Take $b\in X_{+}$. Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1. If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$. Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
Since ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$, there exist $u\in Y_{+}$ and an increasing sequence $(k_{n},z_{n})$ in ${\cal I}$ satisfying ±$\pm T\big{(}\frac{z}{k}\big{)}=\pm Tx_{(k,z)}\leq\frac{1}{n}u$ for all $T\in{\cal T}$ and $(k,z)\geq(k_{n},z_{n})$. In particular, ±$\pm T\big{(}\frac{z}{k_{1}+1}\big{)}=\pm Tx_{(k_{1}+1,z)}\leq u$ for all $T\in{\cal T}$ and $z\in[0,b]$. So, ${\cal T}[0,b]\subseteq[-(k_{1}+1)u,(k_{1}+1)u]$. It follows ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
$ii)$   Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$. For some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $nx_{\alpha}\in[-u,u]$ for all $\alpha\geq\alpha_{n}$. Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$ then ${\cal T}[-u,u]\subseteq[-w,w]$ for some $w\in Y_{+}$, and hence $n{\cal T}x_{\alpha}\subseteq[-w,w]$ for all $\alpha\geq\alpha_{n}$. It follows ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$. ∎
Corollary 2.5.
Order bounded operators between ordered vector spaces with generating cones agree with ru-continuous operators.
Recall that order intervals in an ordered Banach space with a closed cone are norm bounded if and only if the cone is normal (cf., [3, Theorem 2.40]). We include the following elementary collective extension of this fact.
Proposition 2.6.
Let $X$ be an ordered vector space and $Y$ an ordered Banach space with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
Proof.
Let ${\cal T}\in\text{\bf L}_{aob}(X,Y)$ and $[a,b]\subseteq X$. Then ${\cal T}[a,b]\subseteq[-y,y]+B_{Y}$ for some $y\in Y_{+}$. Since $Y_{+}$ is normal, $[-y,y]+B_{Y}$ is norm bounded by [3, Theorem 2.40]. Therefore, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. ∎
The normality of positive cone in the range is essential in Proposition 2.6.
Example 2.7.
Let $X=C^{1}[0,1]$ be an ordered Banach space of continuously differentiable functions on $[0,1]$. The identity operator $I_{X}$ is norm and order bounded yet not order to norm bounded. In particular, $I_{X}\in\text{\bf L}_{aob}(X,Y)\setminus\text{\bf L}_{onb}(X,Y)$.
The following result generalizes [10, Theorem 2.1] to ordered Banach spaces.
Theorem 2.8.
Let $X$ be an ordered Banach space with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
Proof.
Let ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. Suppose, in contrary, ${\cal T}B_{X}$ is not norm bounded. The Krein – Smulian theorem (cf., [3, Theorem 2.37]) implies $\alpha B_{X}\subseteq B_{X}\cap X_{+}-B_{X}\cap X_{+}$ for some $\alpha>0$, and hence ${\cal T}(B_{X}\cap X_{+})$ is not norm bounded. Pick sequences $(x_{n})$ in $B_{X}\cap X_{+}$ and $(T_{n})$ in ${\cal T}$ with $\|T_{n}x_{n}\|\geq n^{3}$ for all $n$, and set $x:=\|\cdot\|$-$\sum\limits_{n=1}^{\infty}n^{-2}x_{n}\in X_{+}$. Since ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ then ${\cal T}[0,x]\subseteq MB_{Y}$ for some $M<\infty$. It follows from $n^{-2}x_{n}\in[0,x]$ that $\|T_{n}(n^{-2}x_{n})\|\leq M$ for every $n$. This is absurd, because $\|T_{n}(n^{-2}x_{n})\|\geq n$ for all $n$. The proof is complete. ∎
Corollary 2.9.
Every order to norm bounded operator from an ordered Banach space $X$ with a closed generating cone to a normed space $Y$ is continuous.
It is worth noting that Corollary 2.9 generalizes the well known fact that every order bounded operator between Banach lattices is bounded. It is also known that every positive operator $T:X\to Y$ is continuous whenever $X$ is an ordered Banach space with a closed generating cone and $Y$ is an ordered Banach space with a closed cone, see Arendt [6, Appendix], Aliprantis and Tourky [3, Theorem 2.32], or Arendt and Nittka [7, Theorem 2.8].
Note that a positive operator need not to be order to norm bounded. Indeed, the identity operator $I_{X}$ on an ordered Banach space $X$ is order to norm bounded if and only if the positive cone $X_{+}$ is normal. Therefore, the next proposition does not follow from Corollary 2.9. Its proof below is a modification of the proof of [3, Theorem 2.32].
Proposition 2.10.
Let $X$ and $Y$ be ordered Banach spaces with closed generating cones. Then every order bounded operator $T:X\to Y$ is continuous.
Proof.
Let $T:X\to Y$ be order bounded. To establish that $T$ is continuous, it suffices to show that $T$ has a closed graph. So, assume $x_{n}\operatorname{\xrightarrow[]{\|\cdot\|}}0$ in $X$ and $Tx_{n}\operatorname{\xrightarrow[]{\|\cdot\|}}y$ in $Y$. We must show that $y=0$.
According to [3, Lemma 2.30], there exist a subsequence $(n_{k})$ such that $x_{n_{k}}\operatorname{\xrightarrow[]{ru}}0$. By Corollary 2.5, $T$ is ru-continuous, and hence $Tx_{n_{k}}\operatorname{\xrightarrow[]{ru}}0$. Thus, there exist some $u\in Y$ and a further subsequence $(n_{k_{m}})$ such that ±$\pm Tx_{n_{k_{m}}}\leq\frac{1}{m}u$, and hence $\frac{1}{m}u\pm Tx_{n_{k_{m}}}\in Y_{+}$ for each $m$. Since $Y_{+}$ is closed and $\frac{1}{m}u\pm Tx_{n_{k_{m}}}\operatorname{\xrightarrow[]{\|\cdot\|}}\pm y$, we conclude that ±$\pm y\in Y_{+}$. That is, y = 0 and the proof is finished. ∎
We have the following collective extension of Proposition 2.10.
Proposition 2.11.
Let $X$ and $Y$ be ordered Banach spaces with closed generating cones, and $X^{\sim}\neq\{0\}$. The following assertions are equivalent.
  1. $i)$
    Every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is norm bounded.
  2. $ii)$
    The cone $Y_{+}$ is normal.
Proof.
$i)\Longrightarrow ii)$ Assume that $Y_{+}$ is not normal. Then, there exists $y_{0}\in X_{+}$ such that the interval $[0,y_{0}]$ is not bounded. Pick $f_{0}\neq 0$ in $X^{\sim}$ and take $x_{0}\in X$ with $f_{0}(x_{0})=1$. Then, $\{f_{0}\otimes y\}_{y\in[0,y_{0}]}\in\text{\bf L}_{ob}(X,Y)$. However, the set $\{f_{0}\otimes y\}_{y\in[0,y_{0}]}$ is not norm bounded because of $\bigcup\limits_{y\in[0,y_{0}]}(f_{0}\otimes y)(x_{0})=[0,y_{0}]$.
$ii)\Longrightarrow i)$ Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$. Then, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$. So, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ by Proposition 2.6. It follows from Theorem 2.8 that ${\cal T}$ is norm bounded. ∎
The following example shows that the condition that $X_{+}$ is closed and generating is essential in Theorem 2.8, in Corollary 2.9, and in Proposition 2.10.
Example 2.12.
Consider an infinite dimensional Banach space $X$.
  1. $a)$
    Let $X_{+}:=\{0\}$ and take an arbitrary discontinuous operator $T$ on $X$. Then $T:(X,X_{+})\to X$ is order to norm bounded and order bounded simply because order intervals in $X$ are singletons. Clearly $X_{+}$ is closed but not generating.
  2. $b)$
    Pick a normalized Hamel basis $\{\xi:\xi\in\Xi\}$ of $X$, identify $X$ with $c_{00}(\Xi)$, and let $X_{+}:=\Big{\{}\sum\limits_{\xi\in\Xi}\lambda_{\xi}\xi:\lambda_{\xi}\geq 0\Big{\}}$. Then $X_{+}-X_{+}=X$ but $X_{+}$ is not closed ((((e.g., by [7, Proposition 4.1])))). Pick any sequence $(\xi_{n})_{n=0}^{\infty}$ of distinct elements of $\Xi$ and define $T:(X,X_{+})\to X$ by $T\xi=0$ for $\xi\in\Xi\setminus\{\xi_{n}:n\in\mathbb{N}\}$ and $T\xi_{n}=n\xi_{0}$ for $n\in\mathbb{N}$. Clearly, $T$ is order to norm bounded and positive, yet not continuous.
Example 2.13.
Let $X$ be an ordered vector space with a generating cone which satisfies the dominated decomposition property ((((i.e., $[0,v_{1}]+[0,v_{2}]=[0,v_{1}+v_{2}]$ for every $v_{1},v_{2}$ in $X_{+}$)))), and let $Y$ be a Dedekind complete vector lattice. It is well known that the ordered vector space ${\cal L}_{ob}(X,Y)$ of order bounded operators of ${\cal L}(X,Y)$ is a Dedekind complete vector lattice (cf., [19, Theorem 83.4]). Let $A$ be an order bounded subset of ${\cal L}_{ob}(X,Y)$, say $A\subseteq[-T,T]$ for some positive operator $T\in{\cal L}_{ob}(X,Y)$. Let $x\in X_{+}$. Since $|Sz|\leq|S|z\leq Tz\leq Tx$ for all $z\in[0,x]$ and $S\in A$, then the set $A$ is collectively order bounded.
From the other hand, not every collectively order bounded set of operators is order bounded. Indeed, consider $X=\ell^{\infty}$, $Y=\mathbb{R}$, and $A=\{\delta^{n}:n\in\mathbb{N}\}\subset\ell^{\infty}_{+}$, where $\delta^{n}(\mathbb{x})=x_{n}$. The set $A$ is collectively order bounded, yet there is no $f\in\ell^{\infty}_{+}$ with $A\subseteq[0,f]$, since otherwise $f(1,1,\ldots,1,\ldots)\geq\sum_{k=1}^{n}\delta^{k}(1,1,\ldots,1,\ldots)=n$ for every $n\in\mathbb{N}$ that is absurd.
3 Some applications to operator semigroups
Under a semigroup $(T_{s})_{s\geq 0}$ on a vector space $X$ we understand a family of operators on $X$ satisfying $T_{s+t}=T_{s}T_{t}$ for all $s,t\in\mathbb{R}_{+}$, and $T_{0}=I_{X}$, where $I_{X}$ is the identity operator on $X$. We shall also use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on $X$, where $J=\mathbb{R}_{+},\mathbb{R},\mathbb{Z}_{+}$, or $\mathbb{Z}$. Such a semigroup is always commutative. More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [17, p.75]).
Proposition 3.1.
Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an ordered Banach space $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
Proof.
By Theorem 2.8, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$, and hence for all $S\in{\text{\rm co}}(\mathcal{G})$. Thus,
$\bigcup_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}SB_{X}=S_{0}\Big{(}\bigcup_{S\in{\text{\rm co}}(\mathcal{G})}SB_{X}\Big{)}\subseteq S_{0}(MB_{X})=MS_{0}B_{X}.$
Since $S_{0}B_{X}$ is relatively (weakly) compact, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. ∎
Theorem 3.2.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an ordered Banach space $X$ with a closed generating cone, and assume that every order interval in $X$ is weakly compact. Then the net $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
Proof.
Since order intervals in $X$ are weakly compact then every almost order bounded subset of $X$ is relatively weakly compact due to the Grothendieck result (cf., [2, Theorem 3.44]). In particular, $\mathcal{G}$ be a collectively order to norm bounded. By Theorem 2.8, $\|S\|\leq M$ for some $M<\infty$ and all $S\in\mathcal{G}$.
Let $x\in X_{+}$. Then, the set $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is almost order bounded, and hence relatively weakly compact. Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weak cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [17, Theorem 1.5]). Since $x\in X_{+}$ is arbitrary and $X_{+}$ is generating, $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
Corollary 3.3.
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or one-parameter discrete operator semigroup on an order continuous Banach lattice $E$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
The following example shows that the order continuity of the norm in $E$ is essential in Corollary 3.3.
Example 3.4.
(cf., [9, Exercise 2.2.22]) Consider the operator $T:c\to c$ such that $T(a_{1},a_{2},\ldots a_{k},\ldots)=\big{(}\frac{1}{2}a_{1},\frac{2}{3}a_{2},\ldots,\frac{k}{k+1}a_{k},\ldots\big{)}$. Clearly, $\big{\{}T^{n}\big{\}}_{n=1}^{\infty}\in\text{\bf L}_{ob}(c)$ for some $\varepsilon>0$. However, the Cesàro averages $\frac{1}{n}\sum_{j=0}^{n-1}T^{j}$ do not converge in the norm at $(1,1,\ldots 1,\ldots)\in c$.
The following corollary of Theorem 3.2 is immediate, because order intervals in the predual of a Neumann algebra are weakly compact by the Akemann theorem (cf., [9, Theorem 3.3.3]).
Corollary 3.5.
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or one-parameter discrete operator semigroup on the predual $\mathcal{M}_{*}$ of a von Neumann algebra $\mathcal{M}$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on ordered vector spaces.
Definition 3.6.
A semigroup $(T_{s})_{s\geq 0}$ on an ordered vector space $X$ is
  1. $a)$
    locally collectively order bounded ((((shortly, lco-bounded))))
    whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M>0$.
  2. $b)$
    lco-continuous whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M>0$.
  3. $c)$
    lcru-continuous whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ruc}(X)$ for every $M>0$.
A semigroup $(T_{s})_{s\geq 0}$ on a normed ordered vector space $X$ is
  1. $d)$
    lcon-bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$ for every $M>0$.
In several important cases, the condition “for every $M>0$” can be replaced by “for some $\varepsilon>0$” in each of the four items of Definition 3.6.
Proposition 3.7.
Let ${\cal T}=(T_{s})_{s\geq 0}$ be a semigroup on an ordered vector space $X$ such that $X_{+}-X_{+}=X$.
  1. $i)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lco-bounded.
  2. $ii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ruc}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lcru-continuous.
  3. $iii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lco-continuous.
  4. $iv)$
    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lcon-bounded.
Proof.
$i)$ Let $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Let $n\varepsilon\geq M$, and take arbitrary $a\leq b$ in $X$. It follows from $X_{+}-X_{+}=X$ that $[a,b]\subseteq[-a_{0},a_{0}]$ for some $a_{0}\in X_{+}$. Since $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ and $X_{+}$ is generating, $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-a_{1},a_{1}]$ for some $a_{1}\in X$. Similarly, for some $a_{2}\in X$, we have $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}]$ and
$\bigcup\limits_{\varepsilon\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]=T_{\varepsilon}\left(\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\right)\subseteq T_{\varepsilon}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}].$
It follows $\bigcup\limits_{0\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-(a_{1}+a_{2}),a_{1}+a_{2}]$. Repeating n-times the argument, we obtain $a_{1},a_{2},...,a_{n}\in X$ with
$\bigcup\limits_{0\leq s\leq M}T_{s}[a,b]\subseteq\bigcup\limits_{0\leq s\leq n\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq\left[-\sum_{i=1}^{n}a_{i},\sum_{i=1}^{n}a_{i}\right].$ (1)1( 1 )
Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order bounded.
$ii)$ By Theorem 2.4, it follows from $i)$.
$iii)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, and fix $M\geq 0$. Let $n\varepsilon\geq M$ and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Pick a net $g^{(1)}_{{\beta_{(1)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(1)}$ there is an $\alpha_{\beta_{(1)}}$ with ±$\pm T_{s}x_{\alpha}\leq g^{(1)}_{{\beta_{(1)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(1)}}$. Since $T_{\varepsilon}$ is order continuous, $T_{\varepsilon}x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. It follows from $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ that there exists a net $g^{(2)}_{{\beta_{(2)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(2)}$ there is $\alpha_{\beta_{(2)}}$ with ±$\pm T_{s}T_{\varepsilon}x_{\alpha}\leq g^{(2)}_{{\beta_{(2)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(2)}}$. Repeating n-times, we obtain nets
$g^{(1)}_{{\beta_{(1)}}}\downarrow 0,\ g^{(2)}_{{\beta_{(2)}}}\downarrow 0,\ ...\ g^{(n)}_{{\beta_{(n)}}}\downarrow 0$
so that, for each $\beta_{(i)}$, $i=1,2,...,n$, there is $\alpha_{\beta_{(i)}}$ satisfying
±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ \alpha\geq\alpha_{\beta_{(i)}}).$
Consider the net $g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}=\sum_{i=1}^{n}g^{(i)}_{{\beta_{(i)}}}\downarrow 0$. Take an
$\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\geq\alpha_{\beta_{(1)}},\alpha_{\beta_{(2)}},....,\alpha_{\beta_{(n)}}.$
Then, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ i=1,2,...,n).$
Therefore, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
±$\pm T_{s}x_{\alpha}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq n\varepsilon).$ (2)2( 2 )
Since $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ is arbitrary, (2) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq n\varepsilon}\in\text{\bf L}_{oc}(X)$, and hence $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order continuous.
$iv)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Take $n\in\mathbb{N}$ with $n\varepsilon\geq M$. Let $a\in X_{+}$. By Theorem 2.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$. Then
$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C^{n}\|a\|.$ (3)3( 3 )
Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$. As $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order to norm bounded. ∎
Under rather mild assumptions, every order bounded (o-continuous, ru-continuous, order to norm bounded) operator between ordered Banach spaces gives a rise to lco-bounded (resp., lco-bounded, lco-continuous, lcru-continuous, lcon-bounded) operator semigroup $(T_{t})_{t\geq 0}$.
Example 3.8.
Let $S:X\to Y$ be a bounded operator between Banach spaces. Let $Z=X\oplus Y$, define a bounded operator $A:Z\to Z$ by $A(x,y)=(0,Sx)$, and set $T_{t}=e^{tA}=\|\cdot\|$-$\sum_{k=0}^{\infty}\frac{t^{k}A^{k}}{k!}$ for $t\geq 0$. Then $(T_{t})_{t\geq 0}$ is a uniformly continuous semigroup on the Banach space $Z$ such that $T_{t}=I+tA$ for $t\geq 0$ because $A^{2}=0$.
Suppose $X$ and $Y$ are ordered Banach spaces and $X_{+}-X_{+}=X$.
  1. $(a)$
    Let $Y_{+}$ be normal and $S:X\to Y$ be order bounded. Then $S$ is order to norm bounded since $S$ is order bounded and $Y_{+}$ is normal. Theorem 2.8 implies that $S$, and hence $A$ is bounded. Since $A$ is order bounded, it follows from $T_{t}=I+tA$ for $t\geq 0$ that $(T_{t})_{t\geq 0}$ is lco-bounded.
  2. $(b)$
    Let $Y_{+}$ be normal and $S:X\to Y$ be ru-continuous. By Theorem 2.4, $S$ is order bounded, and hence $S$ is order to norm bounded because $Y_{+}$ is normal. Due to Theorem 2.8, $S$ is bounded. Thus $A$ is bounded. Since $A$ is ru-continuous, $T_{t}=I+tA$ implies $(T_{t})_{t\geq 0}$ is lcru-continuous.
  3. $(c)$
    Suppose additionally that $X$ is Archimedean. Let $Y_{+}$ be normal and $S:X\to Y$ be order continuous. By Theorem 2.1, $S$ is order bounded. Since $Y_{+}$ is normal, $S$ is order to norm bounded. Theorem 2.8 implies boundedness of $S$, and hence of $A$. Since $A$ is order continuous and $T_{t}=I+tA$, we have $(T_{t})_{t\geq 0}$ is lco-continuous.
  4. $(d)$
    Let $S:X\to Y$ be order to norm bounded. Then, by Theorem 2.8, $S$ is bounded, and so is $A$. Since $A$ is order to norm bounded, $T_{t}=I+tA$ implies $(T_{t})_{t\geq 0}$ is lcon-bounded.
In recent works of Glück, Kandić, Kaplin, and Kramar Fijavž [13, 15, 16] relatively uniformly continuous semigroups on vector lattices were deeply studied. Using of collectively qualified semigroups is too rough for getting progress in this direction. We postpone the investigation of relatively continuous semigroups on ordered vector spaces to a forthcoming paper.